OpenStudy (anonymous):

Two cars leave Chicago, one traveling east and the other west. After three hours, they are 297 miles apart. If one car is traveling 5 mph faster than the other, what is the speed of each vehicle?

OpenStudy (anonymous):

Let the car towards west = W Car towards east = E Let E be travelling at X MPH The W will be travelling at x +5 mph Because the cars are going in oppisite directions, their relative velocit is the sum their respective velocities. ( using Newtonian ) relative velocity = x + (x +5) In three hours the distance travelled is 3 * relative velocity i. distance = 3*(x + (x+5)) = 297 => 6x + 15 = 297 => x = (297 -15)/6 = 282/6 = 47 Therefore there exact speed: 47 mph and 52 mph